A merchant bought 125 kg of potatoes at rs 25 per. Kg. if 5 kg of potatoes were bad, at what price per kg  must he sell the good potatoes to make a gross profit of 20% on the total cost?

Grade:8

2 Answers

Susmita
5 years ago
His selling price is
(125×25×120)/100
But he sells only (25-5)=20 kg.
So price per kg is
(125×25×120)/(100×20) Rs/kg
Aishwarya
5 years ago
CP of 125 kg potatoes = 25 × 125 = ₹3125
P he must make = 20% = 100 + 20/100 * 3125 = Rs. 3750 
Good potatoes = 125 - 5 = 120 kg 
Atq, he needs to sell 120 kg of potatoes at Rs. 3750
 
Then, SP of 1 kg = 3750/120 = Rs. 31.25
 
Hope that helps!